40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 24 02:12:05 2014, Jenne, Update, LSC, Locking without TRY TRX_avg_40cts_POPDC_avg_200cts.pngLockloss_PRCL_350Hz.pngLockloss_DARM_20Hz.png
    Reply  Thu Apr 24 11:19:50 2014, Koji, Update, LSC, Locking without TRY 
       Reply  Thu Apr 24 14:00:42 2014, Jenne, Update, LSC, Locking without TRY REFLDC_1sec.pngREFLDC_3peaks.pngREFLDC_1peak.pngREFLDC_1peak_zoom.png
          Reply  Mon Apr 28 13:18:18 2014, Jenne, Update, LSC, LSC offsets script modified 
Message ID: 9846     Entry time: Thu Apr 24 02:12:05 2014     Reply to this: 9847
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Locking without TRY 

I tried some locking anyway tonight, even though we don't have TRY. 

The biggest conclusion is that I miss the auto-resonance-finding.  I've been roughly scanning the Y-ALS offset to find the POY zero crossing when I see the resonance on the test mass face cameras. 

The next-biggest conclusion, is that I can hold the PRFPMI close to resonance, using ALS for CARM and DARM.  I was trying to transition DARM to AS55, but I couldn't get the last bit of the way.  That is, I couldn't turn off the ALS control.  So, I think that AS55 wasn't actually holding DARM, until maybe the last moment or so.

Anyhow, here are some time series.  My average TRX value is around 40 counts, and POPDC is maybe 250 counts (just PRMI, POPDC is about 75 counts).  Obviously this is noisy as hell, but I'm not using any IR signals for the arms.  Near the end of this first time series, I am trying to switch to AS55 for DARM.

TRX_avg_40cts_POPDC_avg_200cts.png

Zooming in, my real lockloss is due to PRCL oscillating at ~350 Hz:

Lockloss_PRCL_350Hz.png

However, I also saw ~25Hz peaks in CARM and DARM on the spectra starting to show up, and I see a ~25 Hz oscillation in DARM a few moments after the PRCL lockloss.  (Plot #2 is a zoom of the ~1.1 second mark on Plot #3.)

Lockloss_DARM_20Hz.png


The locking parameters:

CARM:

Input:  Using the new CESAR matrix, -1*ALSX, +1*ALSY.  Beatnotes both move up in freq if temp sliders move up.

Servo: gain = 6, FMs 1, 2, 3, 5, 6, 7, 9 on.  Offset = 0 counts. 

Output = -1*MC2

DARM:

Input:  +1*ALSX, +1*ALSY

Servo:  gain = 4.  FMs 1, 2, 3, 5, 6, 7, 9 on.  Offset = 0 counts.

Output = -1*ETMX, +1*ETMY

PRCL:

Input:  +1*REFL33_I, Norm = +0.01*POPDC, sqrt engaged.

Servo:  acquisition easier with -0.04 or -0.06, less gain peaking at -0.02   FMs 4, 5 on; 2, 3, 6, 9 triggered with 0.5 sec delay.  Servo trigger = POPDC, up 100, down 10.  FM trigger = POPDC, up 300, down 20.

Output = +1*PRM

PRCL ASC off, PRM oplev on.

MICH:

Input:  +1*REFL33_Q, Norm = +0.01*POPDC, sqrt engaged.

Servo:  gain = 2, FMs 4, 5 on; 2, 3 triggered with 0.2 sec delay.  Servo trigger = POPDC, up 100, down 10.  FM trigger = POPDC, up 300, down 20.

Output = +0.5*BS, -0.2625*PRM

PDs:

REFL33 analog gain set to 30 dB for both I&Q.

AS55 set to 0 dB for both I&Q.  AS55 had DC normalization of 80 counts (which was the measured number for PRFPMI when TRX was about 0.1 count this evening)

 

ELOG V3.1.3-